# A Systematic Literature Review on Detecting Software Vulnerabilities with Large Language Models

### v1

@@ 4.3.2. Hybrid
Hybrid architectures combine LLMs with other deep learning architectures to leverage complementary strengths. Often, the LLM is used to produce embeddings that are subsequently processed by neural networks such as Recurrent Neural Networks (RNNs), Convolutional Neural Networks (CNNs), or Graph Neural Networks (GNNs).
RNN
@@
RNNs are neural architectures designed specifically to handle sequential data by modeling temporal and contextual dependencies between elements, such as syntactic, semantic, and control-flow relationships (Allamanis et al., 2018). Prominent RNN variants include Long Short-Term Memory networks (LSTMs) and their bidirectional forms (BiLSTMs), which are designed to handle long-range dependencies effectively. In several studies, BERT embeddings are directly fed into (Bi-)LSTM layers (Uddin et al., 2022; Ziems and Wu, 2021). Other studies enhance this setup by passing code embeddings to (Bi)LSTM modules with added attention layers (Li et al., 2021; Xiong and Dong, 2024).
CNN
@@
CNNs are neural architectures specialized in identifying local spatial patterns or features within data through convolutional filters. Initially developed for image recognition tasks, CNNs have been successfully adapted to textual and sequential data analysis by treating sequences as one-dimensional spatial inputs. This capability allows CNNs to effectively capture localized syntactic and semantic patterns in code sequences, making them useful for tasks such as vulnerability detection (Lin et al., 2020). Some studies apply CNNs directly to code or graph embeddings (Safdar et al., 2023; Zhang et al., 2023a). Others utilize pre-trained embedding weights within TextCNN architectures (Hanif and Maffeis, 2022), or integrate CNNs with attention mechanisms to enhance feature extraction (Wang et al., 2023b).
GNN
GNNs are specialized neural architectures designed to operate directly on graph-structured data, effectively modeling relationships between nodes and their structural dependencies.
Given structured representations of code, such as AST and CFG, several studies combine language models with GNN-based architectures to take advantage of both sequential and structural representations of code (Lin et al., 2020).
A common approach is to pass code embeddings, e.g., generated by CodeBERT or GraphCodeBERT, into GNN-based classifiers (Kong et al., 2024; Liu et al., 2025c; Zheng et al., 2024; Hin et al., 2022; Rangapuram and Ratha, 2025).
Other studies use LLMs to feed graph embeddings (Islam et al., 2023, 2024a; Jiang et al., 2024b) or multiple modalities of a function, including text, graph, and image representations (Ni et al., 2023a) into Graph Convolution Networks (GCNs).
Yang et al. (Yang et al., 2024a) use a GNN as a light-weight adapter layer and concatenate its learned embedding with a fine-tuned LLM.
@@
Some studies combine language models with other architectures that do not rely on RNNs, CNNs, or GNNs. Melo et al. (Melo et al., 2025) evaluate the effectiveness of Sparse Autoencoders (SAEs) when applied to representations from LLMs, which are then classified using traditional machine learning models such as random forests. Peng et al. (Peng et al., 2023) use CodeBERT embeddings as input to a Transformer-based classifier. Xue et al. (Xue et al., 2025) perform feature fusion from multiple pre-trained models and process the combined representation with a Transformer model.
4.4. Adaptation Technique
@@
LLMs are adapted to the vulnerability detection task through various state-of-the-art techniques, which differ in how they modify the model’s parameters. The primary adaptation techniques are prompt engineering, fine-tuning, and pre-training. Each strategy reflects different trade-offs between computational cost, model adaptation, and task-specific performance. In the following, we discuss the adaptation techniques as observed in the surveyed studies in detail.
4.4.1. Prompt Engineering
@@
Prompt engineering refers to the design and structuring of input prompts in a way that elicits desired behaviors or outputs. As LLMs remain sensitive to the phrasing, context, and formatting of prompts, numerous prompting strategies have emerged to improve task alignment and performance.
Zero-Shot Prompting
@@
Zero-shot prompting refers to providing the LLM with a task description and input (i.e., a potentially vulnerable code snippet) without any labeled examples (Ambati et al., 2024; Bae et al., 2024; Bakhshandeh et al., 2023; Chan et al., 2023; Cheshkov et al., 2023; Cipollone et al., 2025; Dai et al., 2025; Dozono et al., 2024; Eberhardt and Milánkovich, 2024; Espinha Gasiba et al., 2024; Ezenwoye et al., 2024; Fu et al., 2023b; Gong et al., 2025; Guo et al., 2024b; Ingemann Tuffveson Jensen et al., 2024; Ji et al., 2024; Khan et al., 2025; Khare et al., 2025; Kholoosi et al., 2024; Li et al., 2025d; Lin and Mohaisen, 2024; Liu et al., 2024f, a; Lu et al., 2024; Luo et al., 2025; Mohajer et al., 2024; Ni et al., 2024a; Noever, 2023; Nong et al., 2024; Ozturk et al., 2023; Pelofske et al., 2024; Purba et al., 2023; Ridley et al., 2023; Saimbhi and Akpinar, 2024; Shimmi et al., 2024b; Steenhoek et al., 2025a; Sultana et al., 2024; Sun et al., 2025; Tamberg and Bahsi, 2025; Tian et al., 2024; Torkamani et al., 2025; Trad and Chehab, 2025; Ullah et al., 2024; Wang et al., 2023a, 2024c; Wen et al., 2024a; Wu et al., 2023; Yıldırım et al., 2024; Yu et al., 2024, 2025; Zhang et al., 2025b; Zhou et al., 2024a, b; Zibaeirad and Vieira, 2025, 2024; Farr et al., 2025; Li et al., 2025a; Liu et al., 2025b; Mou et al., 2025; Munson et al., 2025; Shu et al., 2025; Wang et al., 2025a; Zhang et al., 2024c; Wang et al., 2025b; Xia et al., 2025; Li et al., 2025c). Thus, the model is expected to rely solely on its pre-trained knowledge to perform vulnerability detection. Typical prompt formats include simple binary instructions such as ”Is the following code vulnerable? Yes or No,” or targeted CWE-specific queries like ”Does the following code have a CWE-xxx vulnerability?” or ”Which of these types of vulnerabilities is contained in the code? [CWE-xx, CWE-yy, …]”. Studies often experiment with arbitrarily extensive variations to the base prompt, such as modifying task phrasing (e.g., ”vulnerable”, ”buggy”, ”insecure”), providing additional contextual information (e.g., programming language, CWE descriptions, or code function description), or adopting role-based formulations (e.g., ”Act as a vulnerability detection system” (Liu et al., 2024f)) to guide model behavior. Some studies further extend the basic prompt with reinforcement-like strategies (Bae et al., 2024; Liu et al., 2024c; Dubniczky et al., 2025). For example, Liu et al. (Liu et al., 2024c) implement a scoring system where the model earns points when vulnerabilities are correctly identified and fixed, and is penalized when vulnerabilities remain after the fix. Due to its simplicity, zero-shot prompting is commonly used as a starting point and serves as a baseline for comparison with more advanced prompting techniques.
Few-Shot Prompting / In-Context Learning
@@
Few-shot prompting (Brown et al., 2020), also referred to as in-context learning, extends zero-shot prompting by providing the model with several input-output examples relevant to the task. These examples typically consist of code snippets paired with their labels (e.g., vulnerable or non-vulnerable). By presenting these examples in the prompt, the model is guided to infer the desired task behavior, enabling in-context learning without modifying model weights. The examples used for few-shot prompting can be static, i.e., manually selected and reused across prompts (Li et al., 2024; Szabó and Bilicki, 2023; Ding et al., 2025; Dozono et al., 2024; Ni et al., 2024a; Yin et al., 2024; Ji et al., 2024; Dolcetti et al., 2025; Le et al., 2024; Sovrano et al., 2025; Trad and Chehab, 2025; Yildiz et al., 2025; Mou et al., 2025; Wang et al., 2025b), or vulnerability-specific, i.e., tailored examples targeting particular vulnerability types (Ullah et al., 2024; Farr et al., 2025). In another study, the examples are generated by the LLM itself (Chan et al., 2023). Other common strategies for example selection, as discussed by Steenhoek et al. (Steenhoek et al., 2025a), include using random vulnerable and non-vulnerable examples from the training dataset (Gao et al., 2023a; Zhou et al., 2024b; Mohajer et al., 2024; Zhou et al., 2024a; Ni et al., 2024a; Widyasari et al., 2024; Yang et al., 2025b; Steenhoek et al., 2025a), constructing contrastive pairs, where both a vulnerable code snippet and its corresponding fixed version are provided within the same prompt (Nong et al., 2024; Tamberg and Bahsi, 2025; Ceka et al., 2025; Yildiz et al., 2025; Yang et al., 2025b; Steenhoek et al., 2025a), or selecting examples from the training dataset based on (embedding) similarity to the input (Zhou et al., 2024b; Ni et al., 2024a; Widyasari et al., 2024; Shimmi et al., 2024b; Shu et al., 2025; Steenhoek et al., 2025a).
Retrieval-Augmented Generation (RAG)
@@
RAG (Lewis et al., 2020) extends the few-shot paradigm by dynamically augmenting the prompt with relevant external information retrieved from a private knowledge base. RAG systems use a retriever component to identify and fetch contextually similar entries based on embedding similarity or other relevance metrics. The LLM then works with the provided knowledge to generate the output. A key advantage of RAG over traditional few-shot prompting is its ability to dynamically incorporate domain-specific and up-to-date knowledge, making it especially useful for integrating evolving vulnerability knowledge. In the context of vulnerability detection, RAG-based approaches use a wide range of retrieval content. Some studies retrieve the most similar code snippets or code dependencies along with known vulnerability information (Zhang et al., 2025b; Sheng et al., 2024; Liu et al., 2023a; Yang et al., 2025c; Lu et al., 2024; Wen et al., 2024b; Lingxiang et al., 2025; Zhou et al., 2025b). Others rely on abstract code functionality descriptions to better capture the vulnerability semantics and context of the input code (Shimmi et al., 2024b; Sun et al., 2025; Du et al., 2025; Seo et al., 2025). Additionally, vulnerability reports and CWE descriptions are retrieved to guide the model’s reasoning using structured vulnerability information (Keltek et al., 2024; Kouliaridis et al., 2024; Pelofske et al., 2024; Cao and Jun, 2024; Mathews et al., 2024).
Chain-of-Thought (CoT) Prompting
CoT prompting (Wei et al., 2022) aims at enhancing the reasoning capabilities of LLMs by encouraging them to generate intermediate reasoning steps before arriving at a final prediction. Reasoning behavior is typically triggered by simple cues such as ”Let’s think step by step” (Gao et al., 2023a; Li et al., 2024; Bae et al., 2024; Ding et al., 2025; Pelofske et al., 2024; Zhou et al., 2024a; Ni et al., 2024a; Ji et al., 2024; Sheng et al., 2024; Dolcetti et al., 2025; Yildiz et al., 2025; Ullah et al., 2024; Wen et al., 2025; Li et al., 2025e; Wang et al., 2025b; Huynh et al., 2025), requests for explanation (Mohajer et al., 2024; Cipollone et al., 2025; Yang et al., 2025b), or structured step-by-step instructions (Szabó and Bilicki, 2023; Ingemann Tuffveson Jensen et al., 2024; Ceka et al., 2025; Tamberg and Bahsi, 2025; Du et al., 2024; Liu et al., 2024f; Mao et al., 2025; Sun et al., 2025; Yang et al., 2025c; Yu et al., 2025; Sheng et al., 2024; Zibaeirad and Vieira, 2025; Ullah et al., 2024; Munson et al., 2025; Seo et al., 2025; Wang et al., 2025b). Automated CoT prompt construction has also been explored (Trad and Chehab, 2025).
@@
Several studies adapt and extend CoT prompting with domain-specific techniques. For instance, Nong et al. (Nong et al., 2024) propose a vulnerability-semantics-guided prompt format to guide the model through relevant data and control flow facts. More advanced CoT-inspired prompting techniques include self-consistency (Wang et al., 2023d), where the prompt is run multiple times, and predictions are aggregated (e.g., a file is marked vulnerable if the model returns a positive classification in two out of three runs) (Tamberg and Bahsi, 2025), and Tree-of-Thought (ToT) prompting (Tamberg and Bahsi, 2025; Munson et al., 2025), which branches the reasoning process by generating multiple alternative steps per stage, followed by a selection of the most promising path.
Self-Verification
Self-verification builds on the initial prompt-response interaction by prompting the model to re-evaluate, validate, or refine the output. This technique aims to increase detection accuracy and robustness by introducing an additional reasoning loop, either through self-reflection or through interactions with other components.
In single-model setups, the model is prompted to assess its prior response (self-validation or self-refinement) by verifying the correctness of its output (Li et al., 2024; Xia et al., 2025) or recursively critiquing and improving its initial answer (Tamberg and Bahsi, 2025; Du et al., 2024; Yu et al., 2024). Zibaeirad and Vieira (Zibaeirad and Vieira, 2025) introduce a ”Think & Verify” process: the model first performs initial vulnerability detection and then reassesses its findings, assigning a confidence score and providing a severity rating of the detected issue.
In multi-model settings, self-verification is implemented as an interactive process between components. For example, Wang et al. (Wang et al., 2024c) use a vulnerability description refinement approach, where the output of one model is used to prompt another with corrective cues such as ”An expert has found… please recheck it” (Wang et al., 2024c). Eberhardt and Milánkovich (Eberhardt and Milánkovich, 2024) introduce a framework composed of a controller and an evaluator component that work collaboratively using reflection and clarification questions.
Agentic
Agentic prompting extends traditional prompt engineering by embedding LLMs into autonomous, decision-making loops, following a higher-level behavioral pattern involving multi-step planning, tool use, and self-directed reasoning. Agents may operate individually (Shashwat et al., 2024) or as part of multi-agent systems, where agents specialize in subtasks, interact with tools, and collaboratively refine outputs (Cao and Jun, 2024; Yildiz et al., 2025; Nunez et al., 2024; Guo et al., 2025; Seo et al., 2025; Widyasari et al., 2025; Ahmed et al., 2025). For example, Yildiz et al. (Yildiz et al., 2025) employ Reasoning and Acting (ReAct) agents (Yao et al., 2023), which follow a thought-action-observation cycle, for vulnerability detection in code repositories. Seo et al. (Seo et al., 2025) propose a three-agent framework comprising a similarity analyzer, vulnerability verifier, and code patcher; the similarity analyzer performs semantic and taint analysis on LLM-generated code against a RAG database of recent vulnerabilities, while the verifier retrieves matching entries and constructs a one-shot verification example, explaining how the matched vulnerability manifests and its root cause. Widyasari et al. (Widyasari et al., 2025) present a courtroom-inspired multi-agent framework with four role-specific agents: security researcher, code author, moderator, and review board that collaboratively assess code vulnerabilities through iterative analysis, argumentation, summarization, and a final verdict based on collective reasoning. Ahmed et al. (Ahmed et al., 2025) propose an agent pipeline that sequentially analyzes code via normalization, planning, i.e., summarizing the function and generating a checklist of potential vulnerabilities, identifying required external symbols, vulnerability detection using the prior inputs, and validation, with iterative feedback loops between the detection and validation agents.
4.4.2. Training
@@
The adaptation of LLMs to software vulnerability detection tasks often follows the established pre-training and fine-tuning paradigm. In this context, models are either trained from scratch using large code corpora to capture programming-specific syntax and semantics (Ahmad et al., 2021; Buratti et al., 2020; Ding et al., 2023, 2024b; Hanif and Maffeis, 2022; Huang et al., 2022; Jiang et al., 2024c; Liu et al., 2024e; Pan et al., 2021; Uddin et al., 2022; Xu et al., 2023; Yu et al., 2022; Zhao et al., 2024a), used in their (pre-)trained form without any additional training, freezing the model weights (Islam et al., 2023, 2024a; JianJie and Le, 2023; Le et al., 2025; Ni et al., 2023a; Safdar et al., 2023; Wang et al., 2025c, 2023b; Zhang et al., 2023a; Mahmud et al., 2025; Melo et al., 2025; Hin et al., 2022; Mahmud et al., 2025; Melo et al., 2025; Peng et al., 2023; Rangapuram and Ratha, 2025), or further fine-tuned for domain-specific applications. Fine-tuning using vulnerability-specific datasets is the most widely adopted strategy. Studies adopt a range of fine-tuning strategies that vary in their degree of parameter update and architectural design.
Full-Parameter Fine-Tuning

### v2

@@ 4.3.2. Hybrid (S2)
Hybrid architectures combine LLMs with other learning-based components to leverage complementary strengths. Often, the LLM is used to produce embeddings that are subsequently processed by neural networks such as Recurrent Neural Networks (RNNs), Convolutional Neural Networks (CNNs), or Graph Neural Networks (GNNs).
@@
RNNs (S2.1) are neural architectures designed to handle sequential data by modeling temporal and contextual dependencies between elements, e.g., syntactic, semantic, and control-flow relationships (Allamanis et al., 2018). Prominent RNN variants include Long Short-Term Memory networks (LSTMs) and their bidirectional forms (BiLSTMs), which are designed to handle long-range dependencies effectively. In several studies, BERT embeddings are directly fed into (Bi-)LSTM networks (Pham et al., 2024; Ziems and Wu, 2021; Cheng et al., 2025; Li et al., 2021; Xiong and Dong, 2024).
@@
CNNs (S2.2) are neural architectures specialized in identifying local spatial patterns or features within data through convolutional filters. Initially developed for image recognition tasks, CNNs have been successfully adapted to textual and sequential data analysis by treating sequences as one-dimensional spatial inputs. This capability allows CNNs to effectively capture localized syntactic and semantic patterns in code sequences, making them useful for tasks such as vulnerability detection (Lin et al., 2020). Studies apply CNNs directly to code or graph embeddings (Zhang et al., 2023a; Pham et al., 2024). Others use pre-trained embeddings within TextCNN architectures (Hanif and Maffeis, 2022), or integrate CNNs with attention mechanisms to enhance feature extraction (Wang et al., 2023b; Wu and Zhang, 2024).
GNNs (S2.3) are specialized neural architectures designed to operate directly on graph-structured data, effectively modeling relationships between nodes and their structural dependencies. Given structured representations of code, such as AST and CFG, several studies combine LLMs with GNN-based architectures to take advantage of both sequential and structural representations of code (Lin et al., 2020). A common approach is to pass code embeddings, e.g., generated by CodeBERT or GraphCodeBERT, into GNN-based classifiers (Yu and Zhang, 2025; Tian and Zhang, 2025; Kong et al., 2024; Liu et al., 2025b; Zheng et al., 2024; Hin et al., 2022; Rangapuram and Ratha, 2025; Sun et al., 2025c, 2023; Ngan et al., 2025). Other studies use LLMs to feed graph embeddings (Islam et al., 2023, 2024a; Jiang et al., 2024b; Quan et al., 2025) or multiple modalities of a function, including text, graph, and image representations (Ni et al., 2023a) into Graph Convolution Networks (GCNs). Yang et al. (Yang et al., 2024a) use a GNN as a light-weight adapter layer and concatenate its learned embedding with a fine-tuned LLM. Similarly, Gajjar et al. (Gajjar et al., 2025a) use GCN and GAT to extract structural features, subsequently fusing these graph embeddings with semantic embeddings derived from an LLM.
@@
Some studies (Jiang et al., 2025a; Peng et al., 2023; Xue et al., 2025; Melo et al., 2025; Pham et al., 2024) combine LLMs with Other (S2.4) learning-based components that do not rely on RNNs, CNNs, or GNNs. For example, Melo et al. (Melo et al., 2025) evaluate the effectiveness of Sparse Autoencoders (SAEs) when applied to representations from LLMs, which are then classified using traditional machine learning models such as random forests. Jiang et al. (Jiang et al., 2025a) pass embeddings to a Policy network for prediction. Peng et al. (Peng et al., 2023) use CodeBERT embeddings as input to a Transformer-based classifier. Similarly, Xue et al. (Xue et al., 2025) perform feature fusion from multiple pre-trained models and process the combined representation with a Transformer model.
4.4. Technique (T)
To use LLMs effectively for vulnerability detection, studies employ various techniques ranging from simple prompting to complex system and training designs. We broadly categorize techniques into three distinct strategies: (i) feature extraction, (ii) adaptation, and (iii) orchestration.
4.4.1. Feature Extraction
In the most simple approach, language models are used in their pre-trained form without any additional training to extract features by generating input embeddings (Feature Extraction (T1)) (Jiang et al., 2022; Yu and Zhang, 2025; Tian and Zhang, 2025; Pham et al., 2024; Wu and Zhang, 2024; Haque et al., 2025; Liang et al., 2024; Du et al., 2023; Gujar, 2024; Hin et al., 2022; Islam et al., 2023, 2024a; Jararweh et al., 2025; Jiang et al., 2024b, 2025a; Le et al., 2025; Mahmud et al., 2025; Melo et al., 2025; Ngan et al., 2025; Nguyen et al., 2025c; Ni et al., 2023a; Peng et al., 2023; Rangapuram and Ratha, 2025; Sultan et al., 2024; Sun et al., 2023, 2025c; Wang et al., 2025g, 2023b; Xue et al., 2025; Zhang et al., 2023a; Zhao et al., 2023; Zhu et al., 2025b). These high-dimensional vector representations capture the semantic and syntactic properties of the code and serve as input for downstream classifiers.
4.4.2. Adaptation Technique (T2)
@@
Beyond using LLMs as static feature extractors, recent work increasingly focuses on adapting LLMs to the vulnerability detection task through various state-of-the-art techniques, which differ in how they modify the model’s parameters. The primary adaptation techniques are pre-training, fine-tuning, and prompt engineering. Each technique reflects different trade-offs between computational cost, model adaptation, and task-specific performance. In the following, we discuss these adaptation techniques as observed in the surveyed studies.
Prompt Engineering (T2.1)
@@
Prompt engineering refers to the design and structuring of input prompts in a way that elicits desired behaviors or outputs without requiring any modifications to the model’s weights. There are typically two types of prompts: a system prompt (providing general context or behavioral instructions) and a user prompt (specifying the task and providing the code to be analyzed). As LLMs remain sensitive to the phrasing, context, and formatting of prompts, numerous prompting strategies have emerged to improve task alignment and performance.
@@
In Zero-Shot Prompting (T2.1.1), the model is expected to rely solely on its pre-trained knowledge to perform vulnerability detection. The LLM is provided with the task description and input (i.e., a potentially vulnerable code snippet under detection) (Ambati et al., 2024; Bae et al., 2024; Bakhshandeh et al., 2023; Chan et al., 2023; Cheshkov et al., 2023; Cipollone et al., 2025; Dozono et al., 2024; Dubniczky et al., 2025; Eberhardt and Milánkovich, 2024; Espinha Gasiba et al., 2024; Ezenwoye et al., 2024; Farr et al., 2025; Gong et al., 2025; Guo et al., 2024b; Hairan and Şahman, 2025; Huynh et al., 2025; Ingemann Tuffveson Jensen et al., 2024; Khare et al., 2025; Kholoosi et al., 2024; Kouliaridis et al., 2024; Lin and Mohaisen, 2024; Li et al., 2025j; Liu et al., 2024f, b; Li et al., 2025c; Lu et al., 2024; Luo et al., 2025; Maligazhdarova et al., 2025; Ni et al., 2025; Noever, 2023; Ouchebara and Dupont, 2025; Ozturk et al., 2023; Purba et al., 2023; Ridley et al., 2023; Safdar et al., 2025a; Saimbhi and Akpinar, 2024; Shu et al., 2025; Simoni et al., 2025; Steenhoek et al., 2025b, a; Sun et al., 2025b; Trad and Chehab, 2025; Tsai et al., 2025a; Ullah et al., 2024; Wang et al., 2025a, d, b; Wei et al., 2025; Wu et al., 2023; Yıldırım et al., 2024; Yu et al., 2025; Zhang et al., 2025b; Zhou et al., 2024a, b; Zibaeirad and Vieira, 2025, 2024). Typical prompts include simple binary instructions such as ”Is the following code vulnerable? Yes or No,” targeted CWE-specific queries such as ”Does the following code have a CWE-xxx vulnerability?”, or openly phrased ”Does this code have any vulnerabilities?” Studies often experiment with arbitrary variations to the base prompt, such as modifying task phrasing (e.g., ”vulnerable”, ”buggy”, ”insecure”). In a role-based setting, studies include a role in the prompt (e.g., ”Act as a vulnerability detection system” (Liu et al., 2024f)) to guide model behavior. Due to its simplicity, zero-shot prompting is commonly used as a starting point and serves as a baseline for comparison with more advanced prompting techniques.
In-Context Learning (T2.1.2) refers to the prompt engineering technique of embedding additional contextual information (drawing upon the auxiliary information discussed, see Section 4.2) into the model’s input to improve detection without parameter updates. Researchers frequently provide vulnerability definitions, e.g., a candidate list of CWEs (Dozono et al., 2024; Bakhshandeh et al., 2023; Shashwat et al., 2024; Steenhoek et al., 2025a; Wang et al., 2023a; Xia et al., 2025; Yu et al., 2025; Fu et al., 2023b) or detailed vulnerability descriptions (Cao and Jun, 2024; Shimmi et al., 2024b; Gnieciak and Szandala, 2025; Mathews et al., 2024; Mohajer et al., 2024; Ullah et al., 2024). Adopting a code-centric perspective, studies also enrich prompts with execution context (Sayagh and Ghafari, 2025; Sun et al., 2025b; Tamberg and Bahsi, 2025; Zhang et al., 2025a; Ahmed et al., 2025; Mhatre et al., 2025; Li et al., 2025f) or descriptions of the code’s intended functionality (Liu et al., 2024f, 2023; Yang et al., 2025b; Torkamani et al., 2025), supporting the vulnerability detection task. Other approaches use a hybrid workflow by embedding prior detection results from static analysis or similar tools, asking the LLM to verify or refine these findings (Keltek et al., 2024; Munson et al., 2025; Wang et al., 2025e; Yang et al., 2025c).
@@
Few-Shot Prompting (T2.1.3) (Brown et al., 2020), a specific in-context learning technique, provides the model with one or multiple labeled input-output examples relevant to the task. By presenting these examples in the prompt, the model is guided to infer the desired task behavior, enabling in-context learning without modifying model weights. The examples used for few-shot prompting can be static, i.e., manually selected and reused across prompts (Ding et al., 2025; Dolcetti et al., 2025; Le et al., 2024; Li et al., 2025g; Mao et al., 2025; Ni et al., 2025; Sovrano et al., 2025; Trad and Chehab, 2025; Wang et al., 2025b; Yildiz et al., 2025; Yin et al., 2024; Chen et al., 2025a), or vulnerability-specific, i.e., tailored examples targeting particular vulnerability types (Farr et al., 2025; Ouchebara and Dupont, 2025; Ullah et al., 2024). In another study, the examples are generated by the LLM itself (Chan et al., 2023). Other common strategies for example selection, as discussed by Steenhoek et al. (Steenhoek et al., 2025a), include using random vulnerable and non-vulnerable examples from the training dataset (Li et al., 2025j; Mohajer et al., 2024; Ni et al., 2025; Ouchebara and Dupont, 2025; Steenhoek et al., 2025a; Widyasari et al., 2024; Yang et al., 2025b; Zhou et al., 2024a, b), constructing contrastive pairs where both a vulnerable code snippet and its corresponding fixed version are provided within the same prompt (Ceka et al., 2025; Nong et al., 2024; Steenhoek et al., 2025a; Yang et al., 2025b; Yildiz et al., 2025), or selecting examples from the training dataset based on similarity to the input code (Hannan et al., 2025; Ni et al., 2025; Shu et al., 2025; Steenhoek et al., 2025a; Widyasari et al., 2024; Zhou et al., 2024b; Tsai et al., 2025b).
@@
Retrieval-Augmented Generation (T2.1.4) (Lewis et al., 2020) extends the few-shot paradigm by dynamically augmenting the prompt with relevant external information retrieved from a private knowledge base. Retrieval-Augmented Generation (RAG) systems use a retriever component to identify and fetch contextually similar entries based on embedding similarity or other relevance metrics, e.g., BM25. The LLM then works with the provided knowledge to generate the output. A key advantage of RAG over few-shot prompting is its ability to dynamically incorporate domain-specific and up-to-date knowledge, making it especially useful for integrating evolving vulnerability knowledge. In the context of vulnerability detection, RAG-based approaches use a wide range of retrieval content. Some studies retrieve the most similar code snippets or code dependencies along with known vulnerability information (Liu et al., 2024f, 2023; Lu et al., 2024; Ouchebara and Dupont, 2025; Safdar et al., 2025a; Sheng et al., 2024; Steenhoek et al., 2025a; Tsai et al., 2025a; Wen et al., 2024b; Yang et al., 2025c; Zhang et al., 2025b; Zhou et al., 2025b; Zhu et al., 2025a; Jiao et al., 2025; Tsai et al., 2025b). Others rely on abstract code functionality and vulnerability descriptions to better capture the vulnerability semantics and context of the input code (Du et al., 2025a; Li et al., 2025g; Shimmi et al., 2024b; Sun et al., 2025b). Additionally, vulnerability reports and CWE descriptions are retrieved to guide the model’s reasoning using structured vulnerability information (Cao and Jun, 2024; Keltek et al., 2024; Kouliaridis et al., 2024; Mathews et al., 2024; Pelofske et al., 2024; Sayagh and Ghafari, 2025).
@@
Chain-of-Thought (CoT) Prompting (T2.1.5) (Wei et al., 2022) aims at enhancing the reasoning capabilities of LLMs by encouraging intermediate reasoning steps before arriving at a final prediction. Reasoning behavior is typically triggered by simple cues such as ”Let’s think step by step” (Bae et al., 2024; Ding et al., 2025; Dolcetti et al., 2025; Huynh et al., 2025; Ni et al., 2025; Pelofske et al., 2024; Safdar et al., 2025a; Sheng et al., 2024; Steenhoek et al., 2025a; Ullah et al., 2024; Wang et al., 2025b; Wen et al., 2025; Zhou et al., 2024a; Zibaeirad and Vieira, 2025; Yildiz et al., 2025), requests for explanation (Li et al., 2025c; Yang et al., 2025b; Chen et al., 2025a; Khare et al., 2025), or specific step-by-step instructions (Ceka et al., 2025; Du et al., 2024; Li et al., 2025g, i, h; Liu et al., 2024f; Mao et al., 2025; Peng et al., 2025; Sheng et al., 2024; Simoni et al., 2025; Sun et al., 2025b; Ullah et al., 2024; Wang et al., 2025b; Widyasari et al., 2024; Yang et al., 2025c; Yu et al., 2025; Zhou et al., 2025b; Zhu et al., 2025a; Zibaeirad and Vieira, 2025). Automated CoT prompt construction has also been explored (Trad and Chehab, 2025). Several studies adapt CoT prompting with domain-specific techniques. For instance, Nong et al. (Nong et al., 2024) propose a vulnerability-semantics-guided prompt format to guide the model through relevant data and control flow facts. More advanced CoT-inspired prompting techniques include self-consistency (Wang et al., 2023d), where the prompt is run multiple times, and predictions are aggregated (e.g., a file is marked vulnerable if the model returns a positive classification in two out of three runs) (Tamberg and Bahsi, 2025), and Tree-of-Thought (ToT) prompting (Tamberg and Bahsi, 2025; Munson et al., 2025), which branches the reasoning process by generating multiple alternative steps per stage, followed by a selection of the most promising path.
Training (T2.2)
@@
The adaptation of LLMs to software vulnerability detection tasks often follows the established pre-training and fine-tuning paradigm: models are trained from scratch (Pre-Training (T2.2.1)) using large code corpora to capture programming-specific syntax and semantics (Alrashedy et al., 2024; Ahmad et al., 2021; Buratti et al., 2020; Ding et al., 2023, 2024b; Hanif and Maffeis, 2022; Huang et al., 2022; Jararweh et al., 2025; Jiang et al., 2024c; Li et al., 2025e; Liu et al., 2024e; Wang et al., 2025f; Xu et al., 2023; Yu et al., 2022) and further fine-tuned for domain-specific applications. Fine-Tuning (T2.2.2) is the most widely adopted technique. Studies adopt a range of fine-tuning strategies that vary in their degree of parameter update and architectural design.